# OctFusion: Octree-based Diffusion Models for 3D Shape Generation

### v1

@@ 3.1. Octree-based Latent Representation
(2) $L_{sdf}=\frac{1}{N_{\mathcal{Q}}}\sum_{x\in\mathcal{Q}}\left(\lambda_{s}\|F_{sdf}(x)-D(x)\|_{2}^{2}+\|\nabla F_{sdf}(x)-\nabla D(x)\|_{2}^{2}\right),$
where $D(x)$ and $\nabla D(x)$ are the ground-truth SDF and the corresponding gradient at the sampled point $x$, respectively, and $\lambda_{s}$ is set to 200. The second term in Eq. 2 is used to encourage the predicted SDF to be smooth (Wang et al., 2022). To improve the efficiency of the following diffusion model, we also reduce the depth of the original octree with the VAE encoder. Thus, there is an additional binary cross-entropy loss $L_{octree}$ for the splitting status of each octree node in the decoder following (Wang et al., 2022). Finally, we use a KL-divergence loss $L_{KL}$ to regularize the distribution of latent features (Kingma and Welling, 2013) to be similar to a standard Gaussian distribution.
@@
(3) $L_{VAE}=L_{sdf}+L_{octree}+\lambda L_{KL},$
where $\lambda$ is set to 0.1 to balance the affect of $L_{KL}$. The network of VAE comprises residual blocks built upon dual octree graph networks (Wang et al., 2022), downsampling, and upsampling modules, which are detailed in the supplementary material.
Figure 3. Octree-based Latent Representation. Here 2D figures are used for better visualization. Left: An input point cloud sampled from a 3D shape. Middle: The octree structure constructed from the point cloud and latent features for octree leaf nodes produced by the VAE. The latent features are shown as solid dots. Right: The continuous SDF field reconstructed decoded from the latent features. The field value of an arbitrary query point $p$ is computed with the MPU module.
3.2. Octree-Based Diffusion Model
Diffusion Models
A denoising diffusion model (Kingma et al., 2021; Ho et al., 2020) consists of a forward and a reverse process. The forward process is a fixed Markov chain that transforms the data distribution to a Gaussian distribution $\mathcal{N}(0,\boldsymbol{I})$ by iteratively adding noise with the following formula:
(4) $x_{t}=\sqrt{\alpha(t)}x_{0}+\sqrt{1-\alpha(t)}\epsilon,$
@@ 3.2. Octree-Based Diffusion Model
(5) $L_{diffusion}(x_{0})=\mathbb{E}_{\epsilon,t}\|\mathcal{F}(x_{t},t)-x_{0}\|_{2}^{2}.$
@@
After training, we leverage the trained model to sample a generative result from the standard Gaussian distribution (Ho et al., 2020).
Octree-based Diffusion Model
@@
Our octree-based representation is determined by the splitting status and the latent feature of each octree node. The latent features are continuous signals, and we also regard the splitting status as a 0/1 continuous signal, with 00 indicating no splitting and $1$ indicating splitting. Then, we follow Eq. 4 to add noise to the splitting signal and the latent feature of each octree node to get a noised octree. The goal of our OctFusion is to revert the noising process by predicting clean signals for all octree nodes from a noised octree. To this end, we train an octree-based U-Net (Wang et al., 2022, 2017) by minimizing the loss function defined in Eq. 5. The predicted splitting signals are rounded to 0/1 and used to generate the octree structure, and the predicted latent features are used to reconstruct the continuous fields.
@@
We force the octree to be full when the depth is less than $4$ and denote its maximum depth as $D$. The noise is added to octree nodes from depth $4$ to $D$; thus, we need to denoise for multiple octree levels. Instead of training a separate U-Net for each octree level in the spirit of cascaded diffusion models (Zheng et al., 2023; Ren et al., 2024), we propose a unified U-Net that consists of multiple stages and can take noise signals at different octree levels as input and predict the corresponding clean signals. The detailed architecture of our OctFusion is shown in Fig. 2, where each stage of the U-Net is marked with a different color and is responsible for processing octree nodes at a specific depth. We denote the $i^{th}$ stage of the U-Net as $\mathcal{F}_{i}$. The first stage $\mathcal{F}_{1}$ processes octree nodes with depth $4$, which are equivalent to full voxels with resolution $16^{3}$ since the octree is full at this depth. We store an $8$-channel 0/1 signal for each octree node at depth $4$, with which the octree nodes can be split to depth $6$. Sequentially, $\mathcal{F}_{2}$ processes octree nodes with depth $6$ and generated octree with depth $8$, until the last octree depth $D$ is reached. The last stage $\mathcal{F}_{k}$ processes octree nodes with depth $D$ and predicts clean latent features for all leaf nodes.
@@
Although all stages of the U-Net can be trained jointly, we empirically find that training the U-Net stage by stage can produce quantitatively better results. Specifically, we first train $\mathcal{F}_{1}$ at depth $4$, then fix the weights of $\mathcal{F}_{1}$ and train $\mathcal{F}_{2}$, until all stages of the U-Net are is trained. Notebly, when training $\mathcal{F}_{i}$, we reuse the trained weights of $\mathcal{F}_{i-1}$, which enables parameter sharing across different octree levels and is beneficial for training with limited data. It also simplifies the network architecture and improves the efficiency of the model by avoiding the training of multiple separate U-Nets from scratch, greatly reducing the training cost when the octree depth is large. The U-Net uses similar modules as VAE and uses additional attention modules following (Zheng et al., 2023). The detailed network architecture is provided in the supplementary material.
@@
Figure 4. Inference of OctFusion. Here 2D figures are used for better visualization. Top: The first stage model $\mathcal{F}_{1}$ on full octree with depth 4 is used to generate the splitting signals. Bottom: A dual octree graph convolution network $\mathcal{F}_{k}$ is used to generate latent features for all leaf nodes, which will be further decoded to the continuous fields.
@@
To generate a result, we first sample random noise at depth $4$ and generate the splitting signals with $\mathcal{F}_{1}$, with which the octree is grown to depth $6$; then we sample noise at depth $6$ and grow the octree to depth 8 with $\mathcal{F}_{2}$, and so on. In the last stage, we generate latent features for all leaf nodes with $\mathcal{F}_{k}$, which are further decoded to continuous SDFs with the decoder of VAE. A 2D illustration of the sample pipeline is shown in Fig. 4. We then extract the zero-level set of SDFs as generated meshes with marching cubes.
Textured Shape Generation
@@ 4. Experiments
In this section, we verify the effectiveness and generative quality of OctFusion. The experiments can be conducted with 4 Nivdia GPUs with 24G memory.
4.1. 3D Shape Generation
Dataset
@@
We choose 5 categories from ShapeNetV1 (Chang et al., 2015) following LAS-Diffusion (Zheng et al., 2023) and use the same training, evaluation, and testing data split to train our model and conduct comparisons. The categories include chair, table, airplane, car, and rifle. Some meshes in ShapeNet are non-watertight and non-manifold; we repair them following (Wang et al., 2022) and normalize them to the unit cube. We further convert the repaired meshes to signed distance fields (SDFs) for the VAE’s training.
Training Details
@@
To train the VAE, we sample $200k$ points with oriented normals on each repaired mesh and build an octree with depth 8 (resolution $256^{3}$). In each training iteration, we randomly sample $50k$ SDF values from the 3D volume for each shape to evaluate the loss function 3. We train the VAE with the AdamW (Loshchilov and Hutter, 2019) optimizer for 200 epochs with batch size 8 on 2 Nvidia A40 GPUs. The initial learning rate is set as $10^{-3}$ and decreases to $10^{-5}$ linearly throughout the training process. The encoder of VAE downsamples the input octree to depth 6 (resolution $64^{3}$) and outputs a 3-dimensional latent code per octree leaf node. We also train our diffusion models with the AdamW optimizer. The U-Net of OctFusion contains 2 stages, which are trained for 4000 epochs in less than one day and 500 epochs for two days on 4 Nvidia 4090 GPUs, respectively, with a fixed learning rate of $10^{-4}$. To compare with other methods, we train our geometry-only OctFusion model with both unconditional and category-conditional settings. For the unconditional generation, we train our OctFusion on each category. For category-conditional generation, we train our OctFusion on 5 categories data with the label embedding as conditional input.
@@
Figure 5. Examples of generated airplanes and chairs obtained by OctFusion and other state-of-the-art generation models.
Evaluation Metrics
@@
In line with prior research (Zheng et al., 2023; Zhang et al., 2023), we utilize the shading-image-based FID as the primary metric to assess the quality and diversity of the generated 3D shapes. Specifically, we render the generated meshes from 20 uniformly distributed viewpoints, and these images are used to calculate the FID scores against the rendered images from the original training dataset. A lower FID score indicates better generation quality and diversity. Additionally, we adopt the COV, MMD, and 1-NNA metrics (Hui et al., 2022; Achlioptas et al., 2018; Yang et al., 2019), in which COV measures the coverage of the generated shapes, MMD evaluates the fidelity of the generated shapes, and 1-NNA assesses the diversity of the generated shapes. For these metrics, we generate 2000 shapes and compare them with the test set following SDF-StyleGAN (Zheng et al., 2022) and LAS-Diffusion (Zheng et al., 2023). For each mesh, we uniformly sample 2048 points and normalize them within a unit cube to compute the Chamfer distance (CD) and Earth mover’s distance (EMD). Lower MMD, higher COV, and a 1-NNA value closer to 50% indicate better quality.
@@
Table 1. The quantitative comparison of shading-image-based FID between results generated by OctFusion and other methods. The superscript and denote unconditional and category-conditional version of corresponding methods, respectively. Note that Wavelet-Diffusion was trained on 3 categories only and SPAGHETTI was trained on 2 categories only.
@@
Table 2. Additional quantitative comparison between different models. The units of CD and EMD are $10^{-3}$ and $10^{-2}$, respectively.
@@
We conduct comparisons with representative state-of-the-art generative models, including IM-GAN (Chen and Zhang, 2019), SDF-StyleGAN (Zheng et al., 2022), Wavelet-Diffusion (Hui et al., 2022), 3DILG (Zhang et al., 2022), MeshDiffusion (Liu et al., 2023a), SPAGHETTI (Hertz et al., 2022), LAS-Diffusion (Zheng et al., 2023), 3DShape2VecSet (Zhang et al., 2023) and XCube (Ren et al., 2024). Among these methods, IM-GAN and SDF-StyleGAN are GAN-based methods, and the others are diffusion-based methods; 3DILG and 3DShape2VecSet are category conditional models; the others are unconditional generative models trained on each category separately. It is worth noting that we did not conduct any post-processing on our generated meshes extracted by marching cube algorithm (Lorensen and Cline, 1987), while some other methods such as MeshDiffusion (Liu et al., 2023a) removes isolated meshes of tiny sizes and applies both remeshing and the standard Laplace smoothing on all the generated meshes.
@@
Figure 6. Unconditional generation results. Please zoom in for better inspection of the complex geometry and thin structure.
@@
Figure 7. Detailed comparison of OctFusion with 3DShape2VecSet, LAS-Diffusion and XCube on chair and car category. From which we can see that our method has a greater capability of capturing geometry details such as the fluting of swivel chairs and the wheel hubs of cars.
@@
Table 1 reports the comparison of the shading-image-based FID. OctFusion achieves the best-generating quality on average under both unconditional and category-conditional settings compared to all previous methods, demonstrating its superiority. The margin of improvement is more significant in the categories of rifle and airplane, which contain more complex structures and thin details. The comparison with 3DILG, 3DShape2VecSet, XCube is for reference only, as their training data are not exactly the same as those of the other methods. Table 2 reports the quantitative comparison results of COV, MMD, and 1-NNA metrics with models using the same train/eval/test split as ours because these metrics require to be calculated and compared with test set. From Table 2 we can see that our method achieves the best COV(EMD) and 1-NNA(EMD) among all the competing methods and is comparable with Wavelet-Diffusion (Hui et al., 2022) on MMD. However, the drawbacks of COV, MMD, and 1-NNA have been confirmed by previous work (Zheng et al., 2022). Thus, we primarily focus on comparing shading-image-based FID.
Fig. 5 provides qualitative results of airplanes and chairs generated by different methods. We can see that the results IM-GAN, SDF-StyleGAN, and MeshDiffusion contain severe distortions and artifacts. Although LAS-Diffusion, 3DShape2VecSet make significant progress in geometry quality, they fail to capture high-detailed geometric features due to the limitation of resolution of their 3D representation, such as the propellers of airplanes and the thin slats of chairs. Besides Fig. 5, we also provide a far more comprehensive qualitative comparison with LAS-Diffusion, 3DShape2VecSet and a more recently proposed method: XCube in Fig. 7. From which we can see that similar to LAS-Diffusion, XCube uses hierarchical sparse voxel and reaches the resolution of $512^{3}$ on ShapeNet. However, there are still certain difficulties for XCube in modeling the fluting of swivel chairs and the wheel hubs of cars. What’s more, XCube On the contrary, our OctFusion generates implicit features on each octree node which demonstrates superiority in capturing the fine geometry details of 3D shapes. Fig. 6 shows more high-quality and diverse generative results by our unconditional OctFusion model trained on five categories of ShapeNet separately. And we provide more generative results in the supplementary materials, it is highly recommended to check them to get a comprehensive perception of our OctFusion’s generative capability.
Shape Diversity
We evaluate the diversity of generated shapes to identify whether OctFusion just simply memorizes the training data. We conducted this analysis on chair category by retrieving the most similar one to the given generated 3D shape in the training set by Chamfer distance metric. Fig. 8-top shows the histogram whose x-axis is Chamer distrance (×$\times 10^{3}$) which demonstrates that most generated shapes are difference from the training set. Fig. 8-bottom presents some generated samples associated with the three most similar shapes retrieved from the training set. Comparing our generated shapes with the most similar one, it can be clearly observed that our OctFusion is able to generate novel shapes with high diversity instead of just memorizing all cases in training data.
@@
Figure 8. Top: The histogram on the distribution of Chamfer distances between the generated chairs and the most similar one in training set. The unit of CD is $10^{-3}$. Bottom: The generated shapes (Left) and the 3 nearest shapes (right) retrieved from the training dataset according to their Chamfer distances.
Model generalizability

### v2

@@ 3.2.2 Octree-based Variational Autoencoders
$L_{sdf}=\frac{1}{N_{\mathcal{Q}}}\sum_{x\in\mathcal{Q}}\left(\lambda_{s}\|F_{sdf}(x)-D(x)\|_{2}^{2}+\|\nabla F_{sdf}(x)-\nabla D(x)\|_{2}^{2}\right),$ (2)
where $D(x)$ and $\nabla D(x)$ are the ground-truth SDF and the corresponding gradient at the sampled point $x$, respectively, and $\lambda_{s}$ is set to 200. The second term in Eq. 2 is used to encourage the predicted SDF to be smooth [Wang2022]. To improve the efficiency of the following diffusion model, we also reduce the depth of the original octree with the VAE encoder. Thus, there is an additional binary cross-entropy loss $L_{octree}$ for the splitting status of each octree node in the decoder following [Wang2022]. Finally, we use a KL-divergence loss $L_{KL}$ to regularize the distribution of latent features [Kingma2013] to be similar to a standard Gaussian distribution.
@@
$L_{VAE}=L_{sdf}+L_{octree}+\lambda L_{KL},$ (3)
where $\lambda$ is set to 0.1 to balance the affect of $L_{KL}$. The network of VAE comprises residual blocks built upon dual octree graph networks [Wang2022], downsampling, and upsampling modules, which are detailed in the supplementary material.
3.3 Octree-Based Diffusion Model
3.3.1 Diffusion Models
A denoising diffusion model [Kingma2021, Ho2020] consists of a forward and a reverse process. The forward process is a fixed Markov chain that transforms the data distribution to a Gaussian distribution $\mathcal{N}(0,\boldsymbol{I})$ by iteratively adding noise with the following formula:
$x_{t}=\sqrt{\alpha(t)}x_{0}+\sqrt{1-\alpha(t)}\epsilon,$ (4)
@@ 3.3.1 Diffusion Models
$L_{diffusion}(x_{0})={E}_{\epsilon,t}\|\mathcal{F}(x_{t},t)-x_{0}\|_{2}^{2}.$ (5)
@@
After training, we leverage the trained model to sample a generative result from the standard Gaussian distribution [Ho2020].
3.3.2 Octree-based Diffusion Model
@@
Our octree-based representation is determined by the splitting status and the latent feature of each octree node. The latent features are continuous signals, and we also regard the splitting status as a 0/1 continuous signal, with 00 indicating no splitting and $1$ indicating splitting. Then, we follow Eq. 4 to add noise to the splitting signal and the latent feature of each octree node to get a noised octree. The goal of our OctFusion is to revert the noising process by predicting clean signals for all octree nodes from a noised octree. To this end, we train an octree-based U-Net [Wang2022, Wang2017] by minimizing the loss function defined in Eq.5. The predicted splitting signals are rounded to 0/1 and used to generate the octree structure, and the predicted latent features are used to reconstruct the continuous fields.
@@
We force the octree to be full when the depth is less than $4$ and denote its maximum depth as $D$. The noise is added to octree nodes from depth $4$ to $D$; thus, we need to denoise for multiple octree levels. Instead of training a separate U-Net for each octree level in the spirit of cascaded diffusion models [Zheng2023, Ren2024], we propose a unified U-Net that consists of multiple stages and can take noise signals at different octree levels as input and predict the corresponding clean signals. The detailed architecture of our OctFusion is shown in Fig. 1, where each stage of the U-Net is marked with a different color and is responsible for processing octree nodes at a specific depth. We denote the $i^{th}$ stage of the U-Net as $\mathcal{F}_{i}$. The first stage $\mathcal{F}_{1}$ processes octree nodes with depth $4$, which are equivalent to full voxels with resolution $16^{3}$ since the octree is full at this depth. We store an $8$-channel 0/1 signal for each octree node at depth $4$, with which the octree nodes can be split to depth $6$. Specifically, if all the 0/1 signals in 8 channels are 0, the current octree node will not be split. If at least one channel is 1, the current octree node will be split to 8 child nodes to obtain the next level octree with depth 5. Then, the 0/1 signals of the 8 channels will correspond to the newly split 8 child nodes. If the signal is 0, the corresponding child node will not be split and if it is 1, the corresponding child node will be further split. Therefore, we can finally obtain the octree with depth 6. Sequentially, $\mathcal{F}_{2}$ processes octree nodes with depth $6$ and generated octree with depth $8$, until the last octree depth $D$ is reached. The last stage $\mathcal{F}_{k}$ processes octree nodes with depth $D$ and predicts clean latent features for all leaf nodes.
@@
Although all stages of the U-Net can be trained jointly, we empirically find that training the U-Net stage by stage can produce quantitatively better results. Specifically, we first train $\mathcal{F}_{1}$ at depth $4$, then fix the weights of $\mathcal{F}_{1}$ and train $\mathcal{F}_{2}$, until all stages of the U-Net are is trained. Notebly, when training $\mathcal{F}_{i}$, we reuse the trained weights of $\mathcal{F}_{i-1}$, which enables parameter sharing across different octree levels and is beneficial for training with limited data. It also simplifies the network architecture and improves the efficiency of the model by avoiding the training of multiple separate U-Nets from scratch, greatly reducing the training cost when the octree depth is large. The U-Net uses similar modules as VAE and uses additional attention modules following [Zheng2023]. The detailed network architecture is provided in the supplementary material.
@@
Figure 3: Inference of OctFusion. Here 2D figures are used for better visualization. Top: The first stage model $\mathcal{F}_{1}$ on full octree with depth 4 is used to generate the splitting signals. Bottom: A dual octree graph convolution network $\mathcal{F}_{k}$ is used to generate latent features for all leaf nodes, which will be further decoded to the continuous fields.
@@
To generate a result, we first sample random noise at depth $4$ and generate the splitting signals with $\mathcal{F}_{1}$, with which the octree is grown to depth $6$; then we generate the splitting signals at depth $6$ and grow the octree to depth 8 with $\mathcal{F}_{2}$, and so on. In the last stage, we generate latent features for all leaf nodes with $\mathcal{F}_{k}$, which are further decoded to continuous SDFs with the decoder of VAE. A 2D illustration of the sample pipeline is shown in Fig. 3. We then extract the zero-level set of SDFs as generated meshes with Marching Cubes [Lorensen1987].
3.3.3 Textured Shape Generation
@@ 4 Experiments
In this section, we verify the effectiveness and generative quality of OctFusion from a variety of respects.
4.1 3D Shape Generation
4.1.1 Dataset
@@
We choose 5 categories from ShapeNetV1 [Chang2015] following LAS-Diffusion [Zheng2023] and use the same training, evaluation, and testing data split to train our model and conduct comparisons. The categories include chair, table, airplane, car, and rifle. Some meshes in ShapeNet are non-watertight and non-manifold; we repair them following [Wang2022] and normalize them to the unit cube. We further convert the repaired meshes to signed distance fields (SDFs) for the VAE’s training.
4.1.2 Training Details
@@
To train the VAE, we sample $200k$ points with oriented normals on each repaired mesh and build an octree with depth 8 (resolution $256^{3}$). In each training iteration, we randomly sample $50k$ SDF values from the 3D volume for each shape to evaluate the loss function 3. We train the VAE with the AdamW [Loshchilov2017] optimizer for 200 epochs with batch size 8 on 2 Nvidia A40 (48G) GPUs. The initial learning rate is set as $10^{-3}$ and decreases to $10^{-5}$ linearly throughout the training process. The encoder of VAE downsamples the input octree to depth 6 (resolution $64^{3}$) and outputs a 3-dimensional latent code per octree leaf node. We also train our diffusion models with the AdamW optimizer. The U-Net of OctFusion contains 2 stages, which are trained for 4000 epochs in less than one day and 500 epochs for two days on 4 Nvidia 4090 GPUs, respectively, with a fixed learning rate of $10^{-4}$. To compare with other methods, we train our geometry-only OctFusion model with both unconditional and category-conditional settings. For the unconditional generation, we train our OctFusion on each category. For category-conditional generation, we train our OctFusion on 5 categories data with the label embedding as conditional input.
@@
Figure 4: Examples of generated airplanes and chairs obtained by OctFusion and other state-of-the-art generation models.
4.1.3 Evaluation Metrics
@@
In line with prior research [Zheng2023, Zhang2023a], we utilize the shading-image-based FID as the primary metric to assess the quality and diversity of the generated 3D shapes. Specifically, we render the generated meshes from 20 uniformly distributed viewpoints, and these images are used to calculate the FID scores against the rendered images from the original training dataset. A lower FID score indicates better generation quality and diversity. Additionally, we adopt the COV, MMD, and 1-NNA metrics [Hui2022, Achlioptas2018, Yang2019], in which COV measures the coverage of the generated shapes, MMD evaluates the fidelity of the generated shapes, and 1-NNA assesses the diversity of the generated shapes. For these metrics, we generate 2000 shapes and compare them with the test set following SDF-StyleGAN [Zheng2022] and LAS-Diffusion [Zheng2023]. For each mesh, we uniformly sample 2048 points and normalize them within a unit cube to compute the Chamfer distance (CD) and Earth mover’s distance (EMD). Lower MMD, higher COV, and a 1-NNA value closer to 50% indicate better quality.
@@
Table 1: The quantitative comparison of shading-image-based FID between results generated by OctFusion and other methods. The superscript and denote unconditional and category-conditional version of corresponding methods, respectively. Note that Wavelet-Diffusion was trained on 3 categories only and SPAGHETTI was trained on 2 categories only.
@@
Table 2: Additional quantitative comparison between different models. The units of CD and EMD are $10^{-3}$ and $10^{-2}$, respectively.
@@ 4.1.4 Comparisons
We conduct comparisons with representative state-of-the-art generative models, including IM-GAN [Chen2019], SDF-StyleGAN [Zheng2022], Wavelet-Diffusion [Hui2022], 3DILG [Zhang2022a], MeshDiffusion [Liu2023c], SPAGHETTI [Hertz2022], LAS-Diffusion [Zheng2023], 3DShape2VecSet [Zhang2023a] and XCube [Ren2024]. Among these methods, IM-GAN and SDF-StyleGAN are GAN-based methods, and the others are diffusion-based methods; 3DILG and 3DShape2VecSet are category conditional models; the others are unconditional generative models trained on each category separately. It is worth noting that we did not conduct any post-processing on our generated meshes extracted by marching cube algorithm [Lorensen1987], while some other methods such as MeshDiffusion [Liu2023c] removes isolated meshes of tiny sizes and applies both remeshing and the standard Laplace smoothing on all the generated meshes.
@@
Figure 5: Unconditional generation results. Please zoom in for better inspection of the complex geometry and thin structure.
@@
Figure 6: Detailed comparison of OctFusion with 3DShape2VecSet, LAS-Diffusion and XCube on chair and car category. From which we can see that our method has a greater capability of capturing geometry details such as the fluting of swivel chairs and the wheel hubs of cars.
@@
Table 1 reports the comparison of the shading-image-based FID. OctFusion achieves the best-generating quality on average under both unconditional and category-conditional settings compared to all previous methods, demonstrating its superiority. The margin of improvement is more significant in the categories of chair and rifle, which contain more complex structures and thin details. The comparison with 3DILG, 3DShape2VecSet, XCube is for reference only, as their training data are not exactly the same as those of the other methods. Table 2 reports the quantitative comparison results of COV, MMD, and 1-NNA metrics with models using the same train/eval/test split as ours because these metrics require to be calculated and compared with test set. From Table 2 we can see that our method achieves the best COV(EMD) and 1-NNA(EMD) among all the competing methods and is comparable with Wavelet-Diffusion [Hui2022] on MMD. However, the drawbacks of COV, MMD, and 1-NNA have been confirmed by previous work [Zheng2022]. Thus, we primarily focus on comparing shading-image-based FID.
@@
Figure 7: Top: The histogram on the distribution of Chamfer distances between the generated chairs and the most similar one in training set. The unit of CD is $10^{-3}$. Bottom: The generated shapes (Left) and the 3 nearest shapes (right) retrieved from the training dataset according to their Chamfer distances.
Fig. 4 provides qualitative results of airplanes and chairs generated by different methods. We can see that the results IM-GAN, SDF-StyleGAN, and MeshDiffusion contain severe distortions and artifacts. Although LAS-Diffusion, 3DShape2VecSet make significant progress in geometry quality, they fail to capture high-detailed geometric features due to the limitation of resolution of their 3D representation, such as the propellers of airplanes and the thin slats of chairs. Besides Fig. 4, we also provide a far more comprehensive qualitative comparison with LAS-Diffusion, 3DShape2VecSet and a more recently proposed method: XCube in Fig. 6. From which we can see that similar to LAS-Diffusion, XCube uses hierarchical sparse voxel and reaches the resolution of $512^{3}$ on ShapeNet. However, there are still certain difficulties for XCube in modeling the fluting of swivel chairs and the wheel hubs of cars. On the contrary, our OctFusion generates implicit features on each octree node which demonstrates superiority in capturing the fine geometry details of 3D shapes. Fig. 5 shows more high-quality and diverse generative results by our unconditional OctFusion model trained on five categories of ShapeNet separately. And we provide more generative results in the supplementary materials.
4.1.5 Shape Diversity
We evaluate the diversity of generated shapes to identify whether OctFusion just simply memorizes the training data. We conducted this analysis on chair category by retrieving the most similar one to the given generated 3D shape in the training set by Chamfer distance metric. Fig. 7-top shows the histogram whose x-axis is Chamer distrance (×$\times 10^{-3}$) which demonstrates that most generated shapes are difference from the training set. Fig. 7-bottom presents some generated samples associated with the three most similar shapes retrieved from the training set. Comparing our generated shapes with the most similar one, it can be clearly observed that our OctFusion is able to generate novel shapes with high diversity instead of just memorizing all cases in training data.
4.1.6 Model generalizability